# Differentially Private Adaptation of Diffusion Models via Noisy Aggregated Embeddings

### v2

@@ Abstract
We introduce a novel method for adapting diffusion models under differential privacy (DP) constraints, enabling privacy-preserving style and content transfer without fine-tuning model weights. Traditional approaches to private adaptation, such as DP-SGD, incur significant computational and memory overhead when applied to large, complex models. In addition, when adapting to small-scale specialized datasets, DP-SGD incurs large amount of noise that significantly degrades the performance. Our approach instead leverages an embedding-based technique derived from Textual Inversion (TI) and adapted with differentially private mechanisms. We apply TI to Stable Diffusion for style adaptation using two private datasets: a collection of artworks by a single artist and pictograms from the Paris 2024 Olympics. Experimental results show that the TI-based adaptation achieves superior fidelity in style transfer, even under strong privacy guarantees.
Machine Learning, ICML
Figure 1: Overview of DPAgg-TI. We first apply Textual Inversion to extract embeddings for each image in the private dataset. These embeddings are then aggregated with differentially private mechansim, incorporating subsampling to produce a private embedding $u^{\star}$. Finally, images are generated using the corresponding token <$S^{\star}$>.
@@ 1 Introduction
In recent years, diffusion models (Ho et al., 2020; Song et al., 2021b), particularly latent diffusion models (Rombach et al., 2022), have spearheaded high quality text-to-image generation, and have been widely adopted by researchers and the general public alike. Trained on massive datasets like LAION-5B (Schuhmann et al., 2022), these models have developed a broad understanding of visual concepts, enabling new creative and practical applications. Notably, tools such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2023) have been made readily accessible for general use. Building on this foundation, efficient adaptation methods such as parameter efficient fine-tuning (PEFT) methods (Hu et al., 2022; von Platen et al., 2022; Ruiz et al., 2023), guidance based approaches (Ho & Salimans, 2021; Kim et al., 2022; Bansal et al., 2024), and pseudo-word generation (Gal et al., 2023) enable users to leverage this extensive pretraining for customizing models that can specialize on downstream tasks with smaller datasets.
@@
The rapid adoption of diffusion models has raised significant privacy and legal concerns. These models are vulnerable to privacy attacks, such as membership inference (Duan et al., 2023), where attackers determine if a specific data point was used for training, and data extraction (Carlini et al., 2023), which enables reconstruction of training data. This risk is amplified during fine-tuning on smaller, domain-specific datasets, where each record has a greater impact. Additionally, reliance on large datasets scraped without consent raises copyright concerns (Vyas et al., 2023), as diffusion models can reproduce original artworks without credit or compensation. These issues highlight the urgent need for privacy-preserving technologies and clearer ethical and legal guidelines for generative models.
Differential privacy (DP) (Dwork et al., 2006; Dwork, 2006) is a widely adopted approach to addressing these challenges, where controlled noise is added during training or inference to prevent information leakage from individual data points while still enabling the model to learn effectively from the overall dataset. One standard approach for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al., 2016), which modifies traditional SGD by adding noise to clipped gradients.
@@
However, applying DP-SGD to train diffusion models poses several challenges. It introduces significant computational and memory overhead due to per-sample gradient clipping (Hoory et al., 2021), which is essential for bounding gradient sensitivity (Dwork et al., 2006; Abadi et al., 2016). DP-SGD is also incompatible with batch-wise operations like batch normalization, as these link samples and hinder sensitivity analysis. Furthermore, training large models with DP-SGD often leads to substantial performance degradation, particularly under realistic privacy budgets since the required noise scales with the gradient norm. Consequently, existing diffusion models trained with DP-SGD are limited to relatively small-scale images. (Dockhorn et al., 2023; Ghalebikesabi et al., 2023).
As a result, recent research has focused on privacy-preserving strategies for adapting diffusion models without the need for full DP-SGD training. One approach adapts large, publicly pre-trained models to new domains under DP constraints, leveraging their representational strengths while reducing computational and memory costs (Ghalebikesabi et al., 2023). Similarly, PEFT methods like DP-LoRA (Yu et al., 2022) fine-tune a small subset of parameters, enabling efficient adaptation with lower privacy costs. Methods like DP-RDM (Lebensold et al., 2024) avoid direct model updates by using retrieval mechanisms that condition image generation on private data retrieved during inference. However, these alternate approaches often fail to capture the detail of style, underscoring the challenges of balancing privacy, efficiency, and generative performance.
@@
Independent of privacy concerns, Textual Inversion (TI) (Gal et al., 2023) provides an effective method for adapting diffusion models to specific styles or content without modifying the model. Instead, TI learns an external embedding vector that captures the style or content of a target image set, which is then incorporated into text prompts to guide the model’s outputs. A key advantage of TI is its ability to compress a style into a compact vector, reducing computational and memory demands while simplifying the application of privacy-preserving mechanisms, as privacy constraints can be applied directly to embeddings rather than the full model. Additionally, since TI avoids direct model optimization, it remains efficient and compatible with DP constraints on smaller datasets.
@@
In this work, we propose a novel privacy-preserving adaptation method for smaller datasets, leveraging TI to avoid the extensive model updates required by DP-SGD or DP-PEFT approaches. Standard TI inherently compresses the target dataset into a single, low-dimensional vector, providing some obfuscation benefits, but it does not offer formal privacy guarantees. To address this limitation, we introduce a private variant of TI, called Differentially Private Aggregation via Textual Inversion (DPAgg-TI) and summarize it in Figure 1. Our method decouples interactions among samples by learning a separate embedding for each target image, which are then aggregated into a noisy centroid. This approach ensures efficient and secure adaptation to private datasets.
@@
Our experiments demonstrate the effectiveness of DPAgg-TI, showing that TI remains robust in preserving stylistic fidelity even under privacy constraints. Applying our method to a private artwork collection by @eveismyname and Paris 2024 Olympics pictograms (Paris 2024, ), we show that DPAgg-TI captures nuanced stylistic elements while ensuring privacy. We observe a trade-off between privacy (controlled by DP parameter $\varepsilon$) and image quality: lower $\varepsilon$ reduces fidelity but maintains the target style under moderate noise. Subsampling further amplifies privacy by reducing sensitivity to individual data points, mitigating noise impact on image quality. This framework enables privacy-preserving adaptation of diffusion models to new styles and domains while protecting sensitive data.
@@
2 Preliminaries and Related Work
@@ 2.1 Diffusion Models
Diffusion models (Ho et al., 2020; Song et al., 2021b, a; Rombach et al., 2022) leverage an iterative denoising process to generate high-quality images that align with a given conditional input from random noise. In text-to-image generation, this conditional input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text. To convert the text prompt into a suitable conditional format, it is first broken down into discrete tokens, each representing a word or sub-word unit. These tokens are then converted into a sequence of embedding vectors $v_{i}$ that encapsulate the meaning of each token within the model’s semantic space. Next, these embeddings pass through a transformer text encoder, such as CLIP (Radford et al., 2021), outputting a single text-conditional vector $y$ that serves as the conditioning input. This vector $y$ is then incorporated at each denoising step, guiding the model to align the output image with the specific details outlined in the prompt.
@@
Textual Inversion (TI) (Gal et al., 2023) is an adaptation technique that enables personalization using a small dataset of typically 3-5 images. This approach essentially learns a new token that encapsulates the semantic meaning of the training images, allowing the model to associate specific visual features with a custom token.
To achieve this, TI trains a new token embedding, denoted as $u$, representing a placeholder token, denoted as $S$. During training, images are conditioned on phrases such as “A photo of $S$” or “A painting in the style of $S$”. However, unlike the fixed embeddings of typical tokens $v_{i}$, $u$ is a learnable parameter. Let $y_{u}$ denote the text conditioning vector resulting from a prompt containing the token $S$. Through gradient descent, TI minimizes the diffusion model loss given in (1) with respect to $u$, while keeping the diffusion model parameters $\theta$ fixed, iteratively refining this embedding to capture the unique characteristics of the training images. The resulting optimal embedding $u_{*}$ is formalized as follows:
$u_{*}=\arg\min_{u}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}[\|\epsilon-\epsilon_{\theta}(x_{t},y_{u},t)\|^{2}].$ (2)
Hence, $u_{*}$ represents an optimized placeholder token $S_{*}$, which can employed in prompts such as “A photo of $S_{*}$ floating in space” or “A drawing of a capybara in the style of $S_{*}$”, enabling the generation of personalized images that reflect the learned visual characteristics.
2.2 Differential Privacy
In this work, we adopt differential privacy (DP) (Dwork et al., 2006; Dwork, 2006) as our privacy framework. Over the past decade, DP has become the gold standard for privacy protection in both research and industry. It measures the stability of a randomized algorithm with respect to changes in an input instance, thereby quantifying the extent to which an adversary can infer the existence of a specific input based on the algorithm’s output.
Definition 2.1 ((Approximate) Differential Privacy).
@@ 2.2 Differential Privacy
To achieve DP, the Gaussian mechanism is often applied (Dwork et al., 2014; Balle & Wang, 2018), adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$. Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as (Balle & Wang, 2018; Mironov, 2017) to calibrate the noise. (Balle & Wang, 2018), where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$. When the context is clear, we may omit the subscript $f$. This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD (Abadi et al., 2016), which applies Gaussian noise during model updates to achieve DP.
@@
This result is well-known (Steinke (2022, Theorem 29)), with tighter amplification bounds available for Gaussian mechanisms (Mironov, 2017).
2.3 Differentially Private Adaptation of Diffusion Models
Recent advancements in applying DP to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs. Dockhorn et al. (Dockhorn et al., 2023) proposed a Differentially Private Diffusion Model (DPDM) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes using DP-SGD. Another common strategy involves training a model on a large public dataset, followed by differentially private fine-tuning on a private dataset, as explored by Ghalebikesabi et al. (2023). While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the fine-tuning phase (Tramèr et al., 2024).
In response to these limitations, various adaptation techniques have emerged. Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained fine-tuning, as in the VIP approach (Yu et al., 2024), which demonstrates the feasibility of applying DP in later adaptation stages. Other approaches explore differentially private learning of feature representations (Sander et al., 2024), aiming to distill private information into a generalized embedding space while maintaining DP guarantees. Although these adaptations are not yet implemented for diffusion models, they lay essential groundwork for developing secure and efficient privacy-preserving generative models.
@@ 3 Differentially Private Adaptation via Textual Inversion
TI is inherently parameter-efficient and offers certain privacy benefits, as information from an entire dataset of images is compressed into a single token embedding vector. This compression limits the model’s capacity to memorize specific images, making data extraction attacks difficult. However, this privacy is merely heuristic and yet to be proven, so TI may still be vulnerable to privacy attacks such as membership inference. A similar adaptation technique with privacy guarantees may therefore be desirable.
@@
Then, we can aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating the centroid. The purpose of this aggregation is to limit the sensitivity of the final output to each $x^{(i)}$. In order to provide DP guanrantees, we also add isotropic Gaussian noise to the centroid. We can therefore define the resulting embedding vector $u^{*}$ as follows:
$u^{*}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (5)
where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on Balle & Wang (2018, Theorem 1):
@@
The noisy centroid embedding $u^{*}$ can then be used to adapt the downstream image generation process. Similar to regular TI’s $u_{*}$, we can use $u^{*}$ to represent a new placeholder token $S^{*}$ that can be incorporated into prompts for personalized image generation. While $u^{*}$ may not fully solve the TI optimization problem presented in (2), it provides provable privacy guarantees, with only a minimal trade-off in accurately representing the style of the target dataset.
@@
$u^{*}=\frac{1}{m}\sum_{u^{(i)}\in D_{\text{sub}}}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (7)

### v3

@@ Abstract
Personalizing large-scale diffusion models poses serious privacy risks, especially when adapting to small, sensitive datasets. A common approach is to fine-tune the model using differentially private stochastic gradient descent (DP-SGD), but this suffers from severe utility degradation due to the high noise needed for privacy, particularly in the small data regime. We propose an alternative that leverages Textual Inversion (TI), which learns an embedding vector for an image or set of images, to enable adaptation under differential privacy (DP) constraints. Our approach, Differentially Private Aggregation via Textual Inversion (DPAgg-TI), adds calibrated noise to the aggregation of per-image embeddings to ensure formal DP guarantees while preserving high output fidelity. We show that DPAgg-TI outperforms DP-SGD finetuning in both utility and robustness under the same privacy budget, achieving results closely matching the non-private baseline on style adaptation tasks using private artwork from a single artist and Paris 2024 Olympic pictograms. In contrast, DP-SGD fails to generate meaningful outputs in this setting.
Figure 1: We compare our method (DPAgg-TI, top) to a baseline applying DP-SGD to Textual Inversion (bottom), using the prompt “an icon of the Eiffel Tower in the style of the Paris 2024 Olympic Pictograms.” While the baseline learns a single embedding over the dataset, our method privately aggregates per-image embeddings. At privacy budget $\varepsilon=1$, DPAgg-TI preserves visual fidelity much better than the baseline, and closely matches the non-private output (left), demonstrating a superior privacy-utility tradeoff.
Figure 2: Overview of DPAgg-TI. We first apply Textual Inversion to extract embeddings for each image in the private dataset. These embeddings are then aggregated with differentially private mechansim, incorporating subsampling to produce a private embedding $u^{*}_{\text{DP}}$. Finally, images are generated using the corresponding token $S^{*}$>.
@@ 1 Introduction
The rapid adoption of diffusion models  (Ho, Jain, and Abbeel 2020; Song et al. 2021; Rombach et al. 2022) has raised significant privacy and legal concerns. These models are vulnerable to privacy attacks, such as membership inference (Duan et al. 2023), where attackers determine if a specific data point was used for training, and data extraction (Carlini et al. 2023), which enables reconstruction of training data. This risk is amplified during fine-tuning on smaller, domain-specific datasets, where each record has a greater impact. Additionally, reliance on large datasets scraped without consent raises copyright concerns (Vyas, Kakade, and Barak 2023), as diffusion models can reproduce original artworks without credit or compensation. These issues highlight the urgent need for privacy-preserving technologies and clearer ethical and legal guidelines for generative models.
@@
Differential privacy (DP) (Dwork 2006) is a widely adopted framework for addressing these challenges. One standard approach for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al. 2016), which modifies traditional SGD by adding noise to clipped gradients. However, applying DP-SGD to train diffusion models poses several challenges. It introduces significant computational and memory overhead due to per-sample gradient clipping (Hoory et al. 2021), which is essential for bounding gradient sensitivity (Dwork et al. 2006; Abadi et al. 2016). DP-SGD is also incompatible with batch-wise operations like batch normalization, as these link samples and hinder sensitivity analysis. Furthermore, training large models with DP-SGD often leads to substantial performance degradation, particularly under realistic privacy budgets since the required noise scales with the gradient norm. Consequently, existing diffusion models trained with DP-SGD are limited to relatively small-scale images. (Dockhorn et al. 2023; Ghalebikesabi et al. 2023).
@@
Independent of privacy concerns, Textual Inversion (TI) (Gal et al. 2023) provides an effective method for adapting diffusion models to specific styles or content without modifying the model. Instead, TI learns an external embedding vector that captures the style or content of a target image set, which is then incorporated into text prompts to guide the model’s outputs. A key advantage of TI is its ability to compress a style into a compact vector, reducing computational and memory demands while simplifying the application of privacy-preserving mechanisms, as privacy constraints can be applied directly to embeddings rather than the full model. Additionally, since TI avoids direct model optimization, it remains efficient and compatible with DP constraints on smaller datasets.
@@
In this work, we propose a novel privacy-preserving adaptation method for smaller datasets, leveraging TI to avoid the extensive model updates required by DP-SGD. Standard TI does not offer formal privacy guarantees, so to address this limitation, we introduce a private variant of TI, called Differentially Private Aggregation via Textual Inversion (DPAgg-TI) and summarize it in Figure 2. Our method decouples interactions among samples by learning a separate embedding for each target image, which are then aggregated into a noisy centroid. This approach ensures efficient and secure adaptation to private datasets.
@@
Our experiments demonstrate the effectiveness of DPAgg-TI, showing that TI remains robust in preserving stylistic fidelity even under privacy constraints. Applying our method to a private artwork collection by @eveismyname and Paris 2024 Olympics pictograms (Paris 2024 2023), we show that DPAgg-TI captures nuanced stylistic elements while ensuring privacy. We observe a trade-off between privacy (controlled by DP parameter $\varepsilon$) and image quality: lower $\varepsilon$ reduces fidelity but maintains the target style under moderate noise. Subsampling further amplifies privacy by reducing sensitivity to individual data points, mitigating noise impact on image quality. This framework enables privacy-preserving adaptation of diffusion models to new styles and domains while protecting sensitive data.
@@
2 Background and Related Work
@@ 2.1 Diffusion Models
Diffusion models (Ho, Jain, and Abbeel 2020; Song et al. 2021; Song, Meng, and Ermon 2021; Rombach et al. 2022) leverage an iterative denoising process to generate high-quality images that align with a given conditional input from random noise. In text-to-image generation, this conditional input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text. To convert the text prompt into a suitable conditional format, it is first broken down into discrete tokens, each representing a word or sub-word unit. These tokens are then converted into a sequence of embedding vectors $v_{i}$ that encapsulate the meaning of each token within the model’s semantic space. Next, these embeddings pass through a transformer text encoder, such as CLIP (Radford et al. 2021), outputting a single text-conditional vector $y$ that serves as the conditioning input. This vector $y$ is then incorporated at each denoising step, guiding the model to align the output image with the specific details outlined in the prompt.
@@
Textual Inversion (TI) (Gal et al. 2023) is an adaptation technique that enables personalization using a small dataset of typically 3-5 images. This approach essentially learns a new token that encapsulates the semantic meaning of the training images, allowing the model to associate specific visual features with a custom token.
To achieve this, TI trains a new token embedding, denoted as $u$, representing a placeholder token, denoted as $S$. During training, images are conditioned on phrases such as “A photo of SS” or “A painting in the style of SS”. However, unlike the fixed embeddings of typical tokens $v_{i}$, $u$ is a learnable parameter. Let $y_{u}$ denote the text conditioning vector resulting from a prompt containing the token $S$. Through gradient descent, TI minimizes the diffusion model loss given in (1) with respect to $u$, while keeping the diffusion model parameters $\theta$ fixed, iteratively refining this embedding to capture the unique characteristics of the training images. The resulting optimal embedding $u^{*}$ is formalized as follows:
$u^{*}=\arg\min_{u}\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,I),t\sim[T]}[\|\epsilon-\epsilon_{\theta}(x_{t},y_{u},t)\|^{2}].$ (2)
Hence, $u^{*}$ represents an optimized placeholder token $S^{*}$, which can employed in prompts such as “A photo of $S^{*}$ floating in space” or “A drawing of a capybara in the style of $S^{*}$”, enabling the generation of personalized images that reflect the learned visual characteristics.
2.2 Differential Privacy.
In this work, we adopt differential privacy (DP) (Dwork et al. 2006; Dwork 2006) as our privacy framework. Over the past decade, DP has become the gold standard for privacy protection in both research and industry. It measures the stability of a randomized algorithm with respect to changes in an input instance, thereby quantifying the extent to which an adversary can infer the existence of a specific input based on the algorithm’s output.
Definition 1 ((Approximate) Differential Privacy).
@@
To achieve DP, the Gaussian mechanism is often applied (Dwork, Roth et al. 2014; Balle and Wang 2018), adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$. Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as (Balle and Wang 2018; Mironov 2017) to calibrate the noise. (Balle and Wang 2018), where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$. When the context is clear, we may omit the subscript $f$. This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD (Abadi et al. 2016), which applies Gaussian noise during model updates to achieve DP.
@@
This result is well-known (Steinke (2022, Theorem 29)), with tighter amplification bounds available for Gaussian mechanisms (Mironov 2017).
2.3 Private Adaptation of Diffusion Models
Recent advancements in applying DP to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs. Dockhorn et al. Dockhorn et al. (2023) proposed a Differentially Private Diffusion Model (DPDM) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes using DP-SGD. Another common strategy involves training a model on a large public dataset, followed by differentially private fine-tuning on a private dataset, as explored by Ghalebikesabi et al. (2023). While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the fine-tuning phase (Tramèr, Kamath, and Carlini 2024).
In response to these limitations, various adaptation techniques have emerged. Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained fine-tuning, as in the VIP approach (Yu et al. 2024), which demonstrates the feasibility of applying DP in later adaptation stages. Other approaches explore differentially private learning of feature representations (Sander et al. 2024), aiming to distill private information into a generalized embedding space while maintaining DP guarantees. Although these adaptations are not yet implemented for diffusion models, they lay essential groundwork for developing secure and efficient privacy-preserving generative models.
@@ 2.3 Private Adaptation of Diffusion Models
Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “A painting of Taylor Swift in the style of <<@eveismyname>>”, with the embedding <<@eveismyname>> trained using different values of $m$ and $\varepsilon$.
Figure 5: Images generated by Stable Diffusion v1.5 using the prompt “Icon of a dragon in the style of <<Paris 2024 Pictograms>>”, with the embedding <<Paris 2024 Pictograms>> trained using different values of $m$ and $\varepsilon$.
@@ 3 Differentially Private Adaptation via Textual Inversion
Then, we can aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating the centroid. The purpose of this aggregation is to limit the sensitivity of the final output to each $x^{(i)}$. In order to provide DP guarantees, we also add isotropic Gaussian noise to the centroid. We can therefore define the resulting embedding vector $u^{*}_{\text{DP}}$ as follows:
$u^{*}_{\text{DP}}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (5)
where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on Balle and Wang (2018, Theorem 1):
@@
The noisy centroid embedding $u^{*}_{\text{DP}}$ can then be used to adapt the downstream image generation process. Similar to regular TI’s $u^{*}$, we can use $u^{*}_{\text{DP}}$ to represent a new placeholder token $S^{*}$ that can be incorporated into prompts for personalized image generation. While $u^{*}_{\text{DP}}$ may not fully solve the TI optimization problem presented in (2), it provides provable privacy guarantees, with only a minimal trade-off in accurately representing the style of the target dataset.
@@
$u^{*}_{\text{DP}}=\frac{1}{m}\sum_{u^{(i)}\in D_{\text{sub}}}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (7)